# Embracing Large Language Models in Traffic Flow Forecasting

### v1

@@ Abstract
Traffic flow forecasting aims to predict future traffic flows based on the historical traffic conditions and the road network. It is an important problem in intelligent transportation systems, with a plethora of methods been proposed. Existing efforts mainly focus on capturing and utilizing spatio-temporal dependencies to predict future traffic flows. Though promising, they fall short in adapting to test-time environmental changes of traffic conditions. To tackle this challenge, we propose to introduce large language models (LLMs) to help traffic flow forecasting and design a novel method named Large Language Model Enhanced Traffic Flow Predictor (LEAF). LEAF adopts two branches, capturing different spatio-temporal relations using graph and hypergraph structures respectively. The two branches are first pre-trained individually, and during test-time, they yield different predictions. Based on these predictions, a large language model is used to select the most likely result. Then, a ranking loss is applied as the learning objective to enhance the prediction ability of the two branches. Extensive experiments on several datasets demonstrate the effectiveness of the proposed LEAF.
@@ 1 Introduction
Traffic flow forecasting is an integral part of intelligent transportation systems Dimitrakopoulos and Demestichas (2010); Zhang et al. (2011) and smart cities Shahid et al. (2021); Dai et al. (2022). The goal of traffic flow forecasting is to predict future traffic flows using the historical data and the spatial information (i.e. the road network), which has a wide range of applications including traffic signal control Jiang et al. (2021), route planning Liebig et al. (2017), and congestion management Fouladgar et al. (2017).
@@
Despite their success, existing traffic flow forecasting methods have two limitations, which hinder their applications in the real world. (1) Unable to adapt to environmental changes of traffic conditions during test time. Most existing methods make the assumption that test data follow the same (or a very similar) distribution as the training data, which may fail to hold true in real-world scenarios Kim et al. (2024); Chen et al. (2024c), especially for time series data Lu et al. (2022); Gagnon-Audet et al. (2022); Jian et al. (2024). In the real world, traffic condition changes over time due to a variety of factors like special events, the change of weather, or the shift of eras. While it is difficult for existing methods to adequately adapt to all these changes, the assistance of large language models (LLMs) can make a difference, as LLMs have the ability to understand these changes Chang et al. (2024); Minaee et al. (2024); Beniwal et al. (2024). A naive solution is to utilize the generative ability of LLMs to make direct predictions Li et al. (2024b); Ren et al. (2024); Liang et al. (2024), as shown in Figure 1. However, directly generating future traffic flows could be too challenging for language models, as accurate forecasting relies on both historical data and complex spatio-temporal relations. (2) Weak in capturing the rich structure of spatio-temporal relations in traffic data. The traffic network is complicated and the temporal dimension adds another layer of complexity. A large number of prior works focus on capturing the complex spatio-temporal relations, using graph structures Song et al. (2020); Zheng et al. (2023a) or hypergraph structures Wang et al. (2022); Wang and Zhu (2022); Zhao et al. (2023). Graphs capture pair-wise relations, while hypergraphs model non-pair-wise relations. Adopting only one of them is not enough, as the spatio-temporal relations in traffic data is rich by nature. For example, traffic congestion at one vertex affects adjacent vertices (pair-wise relations), whereas road closures affect a large set of vertices (non-pair-wise relations). Modeling the rich structure of spatio-temporal relations is a challenging aspect of predicting future traffic flows.
@@
Towards this end, we propose a novel method termed Large Language Model Enhanced Traffic Flow Predictor (LEAF) for adaptive and structure-perspective traffic flow forecasting. LEAF consists of a predictor and a selector, where the predictor generates options of predictions and the selector chooses the most likely result. To enhance adaptability, we build an LLM-based selector that selects from a range of possible future traffic flows using the discriminative ability of a large language model, as shown in Figure 1. The selection results are used to guide the predictor with a ranking loss. The large language model is good at understanding the changing traffic conditions and is open to further information provided by humans, making it an adaptable predictor. To better capture the rich structures of spatio-temporal relations, we build a dual-branch predictor composed of a graph branch which captures pair-wise relations of spatio-temporal traffic data, and a hypergraph branch, which captures non-pair-wise relations. During test time, the dual-branch traffic flow predictor generates different forecasting results, and subsequently, a set of transformations is applied to obtain a wealth of choices of future traffic flows.
@@
  • We propose an LLM-enhanced traffic flow forecasting framework that introduces large language models in test time to enhance the adaptability of traffic flow forecasting models.
@@
  • Extensive experiments on several benchmark datasets verify the effectiveness of the proposed method.
@@ 2.1 Traffic Flow Forecasting
Traffic flow forecasting is a topic that has been studied for several decades Smith and Demetsky (1997); Sun et al. (2006); Yang et al. (2016); Song et al. (2020); Li et al. (2024b). Early efforts mainly focus traditional models Smith and Demetsky (1997); Asadi et al. (2012). With the success of deep learning, deep neural networks have become the mainstream in this field. One line of research adopt recurrent neural networks (RNNs) Hochreiter (1997) and graph neural networks (GNNs) Kipf and Welling (2016), where the GNNs and RNNs capture the spatial and temporal relations respectively Li et al. (2018); Wang et al. (2020); Chen et al. (2022b).
To jointly model spatial and temporal relations, another line of research utilize GNNs in both dimensions Song et al. (2020); Li and Zhu (2021); Lan et al. (2022); Chen et al. (2024a). As simple graphs only capture pair-wise relations, some works make a step further, introducing hypergraph neural networks (HGNNs) to capture non-pair-wise spatio-temporal relations Luo et al. (2022); Wang and Zhu (2022); Zhao et al. (2023); Wang et al. (2024). In this work, we take the benefits from both sides, with a dual-branch predictor, capturing rich structures of spatio-temporal relations.
Figure 2: The framework of the proposed LEAF. LEAF consists of a dual-branch traffic flow predictor and a large language model based selector. The predictor generates forecasts of future traffic flows through the graph branch and the hypergraph branch. The selector constructs choice set and then selects the best option using (frozen) LLMs. The results of selection is used to supervise the predictor.
@@ 2.2 Large Language Models
In recent years, large language models has drawn increased attention, both within the community of natural language processing Chen et al. (2024b); Yin et al. (2024) and beyond. Efforts have been made to utilize the power of LLMs in other fields, including healthcare Liévin et al. (2024); Van Veen et al. (2024); Labrak et al. (2024), education Milano et al. (2023), legal technology Lai et al. (2024), economics Li et al. (2024a), recommendation Chen et al. (2024b); Bao et al. (2024), and transportation Liu et al. (2024); Guo et al. (2024b); Ren et al. (2024).
Among these applications, traffic flow forecasting is an important one, being the foundation of smart cities Boukerche et al. (2020). The success of LLMs in language has inspired a plethora of works in this task. Some works adopt the architecture of LLMs for building transformer-based traffic flow predictors Cai et al. (2020); Xu et al. (2020); Chen et al. (2022a); Liu et al. (2023); Jiang et al. (2023); Zou et al. (2024). However, they typically require a large amount of data for training.
Another line of research attempts to equip LLMs with the ability to predict future traffic flows based on the history and specific situations Zheng et al. (2023b); Guo et al. (2024a); Li et al. (2024b); Yuan et al. (2024); Han et al. (2024). Although LLMs have shown promising results in understanding time series Yu et al. (2023); Koval et al. (2024); Gruver et al. (2024) or temporal events Xia et al. (2024); Hu et al. (2024), the traffic data involves complex spatio-temporal relations challenging LLMs’ generative ability. In this work, we show that one can instead utilize their discriminative ability to enhance existing traffic flow forecasting models.
@@ 3 Methodology
3.1 Problem Setup and Overview
Problem Setup. We follow the standard setup for traffic flow forecasting Song et al. (2020), where there is a road network denoted as $\mathcal{G}=\langle\mathcal{V},\mathcal{E},\bm{A}\rangle$. $\mathcal{V}$ is the set of $N$ vertices (i.e. the sensors in the city), $\mathcal{E}$ is the set of edges, and $\bm{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix. In this road network, historical traffic flows can be represented as $\bm{X}=(\bm{X}_{1},\bm{X}_{2},\cdots,\bm{X}_{T})\in\mathbb{R}^{T\times N\times F}$, where $T$ is the length of historical observation and $F$ is the dimension of input features. The goal is to predict the future of traffic flows with the length of $T^{\prime}$, denoted as $\bm{X}^{\prime}=(\bm{X}^{\prime}_{1},\bm{X}^{\prime}_{2},\cdots,\bm{X}^{\prime}_{T^{\prime}})\in\mathbb{R}^{T^{\prime}\times N\times F}$.
Framework Overview. To solve the aforementioned problem with the help of large language models, we propose a novel framework termed Large Language Model Enhanced Traffic Flow Predictor (LEAF). The overview of the framework is illustrated in Figure 2. Specifically, to achieve LLM-enhanced prediction (Section 3.4), we design a dual-branch traffic flow predictor (Section 3.2) and an LLM-based selector (Section 3.3). The predictor consists of a graph neural network branch capturing pair-wise spatio-temporal relations, and a hypergraph neural network branch capturing non-pair-wise relations. During training, the predictor is first pretrained using the training data. During test time, we apply transformations to the forecasting results of the predictor to obtain a variety of choices, among which the selector chooses the best one with a frozen LLM. The selection results are then used to fine-tune the dual-branch predictor with a ranking loss during test time so as to improve the adaptability of the model.
@@ 3.2 Dual-branch Traffic Flow Predictor
Previous works on traffic flow forecasting adopt the graph perspective Song et al. (2020); Zheng et al. (2023a) or the hypergraph perspective Wang et al. (2022); Zhao et al. (2023). The graph perspective propagates messages between pairs of nodes, which makes them adept in modeling pair-wise spatio-temporal relations (e.g. an accident affects adjacent locations). On the other hand, the hypergraph perspective propagates messages among groups of nodes, making them proficient in modeling non-pair-wise spatio-temporal relations (e.g. people move from the residential area to the business area). For LEAF, we aim to take the benefits from both sides by constructing a dual-branch predictor and letting the LLM to select.
@@
where $\sigma(\cdot)$ is the activation layer and $\bm{W}_{G}^{(l)}\in\mathbb{R}^{d\times d}$ is the weight matrix. $\widehat{\bm{A}^{ST}}=\bm{D}^{-1/2}\bm{A}^{ST}\bm{D}^{-1/2}$ is the normalized version of the adjacency matrix, where $\bm{D}$ is the degree matrix Kipf and Welling (2016); Song et al. (2020). By adopting graph convolutions in Eq. 3, the information from one spatio-temporal vertex can be propagated to its neighbors as defined in $\mathcal{E}^{ST}$ in Eq. 3.2, and thus this branch models pair-wise spatio-temporal relations.
The Hypergraph Branch. Although the graph branch is adept in capturing pair-wise relations, the complex traffic patterns contain non-pair-wise relations. For example, in the morning rush hours, people move from the residential area (which is a set of vertices in a hyperedge) to the business area (which is another hyperedge). The vertices in one hyperedge share common patterns and the hyperedges affect each other. To model non-pair-wise relations, hypergraphs are adopted. For a hypergraph, its incidence matrix $\bm{I}_{H}\in\mathbb{R}^{NT\times m}$ describes the assignment of $NT$ vertices to $m$ hyperedges. As the incidence matrix is not given as input, we resort to a learnable one with low-rank decomposition Zhao et al. (2023):
@@
$\bm{X}^{(l)}_{H}=\bm{I}_{H}\left(\bm{I}_{H}^{T}\bm{X}^{(l-1)}_{H}+\sigma\left(\bm{W}_{E}\bm{I}_{H}^{T}\bm{X}^{(l-1)}_{H}\right)\right),$ (5)
@@
3.3 Large Language Model Based Selector
In the Section 3.2, we obtain different prediction results from different branches, denoted as $\bm{Y}^{G}$ and $\bm{Y}^{H}$. Most previous works Li et al. (2024b); Ren et al. (2024); Liang et al. (2024) directly generate these predictions, which is challenging since the complex spatio-temporal relations are hard to express in texts. By comparison, the LLM-based selector aims to choose the best prediction, using the internal knowledge of a frozen LLM and the constructed prompts. As the traffic networks are usually large, we break the result $\bm{Y}\in\{\bm{Y}^{G},\bm{Y}^{H}\}\subset\mathbb{R}^{T^{\prime}\times N}$ for individual vertices, i.e. $\bm{Y}=[\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{N}]$, where $\bm{y}_{i}\in\mathbb{R}^{T^{\prime}}$.
Choice Set Construction. In practice, we want to give the LLM-based selector more choices so that it has the potential to make better predictions. Therefore, we introduce several transformations: smoothing, upward trend, downward trend, overestimate, and underestimate. The set of transformations is denoted as $\mathcal{T}$. For vertex $i\in\mathcal{V}$, the choice set is determined as follows:

### v2

@@ Abstract
Traffic flow forecasting aims to predict future traffic flows based on historical traffic conditions and the road network. It is an important problem in intelligent transportation systems, with a plethora of methods being proposed. Existing efforts mainly focus on capturing and utilizing spatio-temporal dependencies to predict future traffic flows. Though promising, they fall short in adapting to test-time environmental changes in traffic conditions. To tackle this challenge, we propose to introduce large language models (LLMs) to help traffic flow forecasting and design a novel method named Large Language Model Enhanced Traffic Flow Predictor (LEAF). LEAF adopts two branches, capturing different spatio-temporal relations using graph and hypergraph structures, respectively. The two branches are first pre-trained individually, and during test time, they yield different predictions. Based on these predictions, a large language model is used to select the most likely result. Then, a ranking loss is applied as the learning objective to enhance the prediction ability of the two branches. Extensive experiments on several datasets demonstrate the effectiveness of LEAF. Our code is available at https://github.com/YushengZhao/LEAF.
@@ 1 Introduction
Traffic flow forecasting is an integral part of intelligent transportation systems Dimitrakopoulos and Demestichas (2010); Zhang et al. (2011) and smart cities Shahid et al. (2021); Dai et al. (2022). The target of traffic flow forecasting is to predict future traffic flows using historical data and spatial information (i.e. the road network), which has a wide range of applications including traffic signal control Jiang et al. (2021), route planning Liebig et al. (2017), and congestion management Fouladgar et al. (2017).
@@
Despite their success, existing traffic flow forecasting methods have two limitations, which hinder their applications in the real world.
Firstly, existing methods are generally unable to adapt to environmental changes of traffic conditions during test time. Most existing methods make the assumption that test data follow the same distribution as the training data, which may fail to hold in the real world Lu et al. (2022); Zhao et al. (2025a, c), especially for time series data Kim et al. (2021); Fan et al. (2023); Chen et al. (2024c); Zhang et al. (2024). Traffic conditions change over time due to a variety of factors like special events, changes in weather, or the shift of eras. Traditional models Song et al. (2020); Li and Zhu (2021), including fusion models Kashinath et al. (2021) have poor generalization ability, suffering from performance degradation under distribution shifts. In contrast, LLMs have strong generalization ability to adapt well under such distribution shifts Chang et al. (2024); Minaee et al. (2024); Beniwal et al. (2024). Moreover, LLMs have strong reasoning ability, which enables them to infer the current environment from the data to benefit forecasting. Due to the high cost of collecting large-scale data, the zero-shot generalization/reasoning ability is especially useful. To utilize LLMs, a naive solution is to use the generative ability of LLMs to make direct predictions Li et al. (2024c); Ren et al. (2024); Liang et al. (2024), as shown in Figure 1 (upper part). However, this is too challenging for language models, as accurate forecasting relies on both historical data and complex spatio-temporal relations.
@@
Secondly, existing methods are weak in capturing the rich structure of spatio-temporal relations in traffic data. The traffic network is complicated and the temporal dimension adds another layer of complexity. Prior works focus on capturing the complex spatio-temporal relations, using graph structures Song et al. (2020); Zheng et al. (2023a) or hypergraph structures Wang et al. (2022); Wang and Zhu (2022); Zhao et al. (2023). Graphs capture pair-wise relations, while hypergraphs model non-pair-wise relations. Adopting only one of them is not enough, as the spatio-temporal relations in traffic data are rich by nature. For example, traffic congestion at one vertex affects adjacent vertices (pair-wise relations), whereas road closures affect a large set of vertices (non-pair-wise relations). Modeling the rich structure of spatio-temporal relations is challenging in predicting future traffic flows.
@@
To that end, this paper proposes a novel method termed Large Language Model Enhanced Traffic Flow Predictor (LEAF) for adaptive and structure-perspective traffic flow forecasting. The core idea of LEAF is to utilize the discriminative ability of LLMs to enhance the task of traffic flow forecasting using a predictor and a selector, where the predictor generates predictions and the selector chooses the most likely result. To enhance adaptability, we build an LLM-based selector that selects from a range of possible future traffic flows using the discriminative ability of an LLM, as shown in Figure 1 (lower part). The selection results are used to guide the predictor with a ranking loss Weinberger and Saul (2009); Sohn (2016), which only requires that the positive candidate (the ones chosen by the LLM) is better than the negative candidates (the ones not chosen). The LLM is good at understanding the changing traffic conditions and is open to further information provided by humans, making it an adaptable predictor. To better capture the rich structures of spatio-temporal relations, we build a dual-branch predictor composed of a graph branch which captures pair-wise relations of spatio-temporal traffic data, and a hypergraph branch, which captures non-pair-wise relations. During test time, the dual-branch predictor generates different forecasting results, and subsequently, a set of transformations is applied to obtain a wealth of choices of future traffic flows.
@@
  • We propose an LLM-enhanced traffic flow forecasting framework that introduces LLMs in test time to enhance the adaptability under distribution shift of traffic flow forecasting models.
@@
  • Extensive experiments on several datasets verify the effectiveness of the proposed method.
@@ 2.1 Traffic Flow Forecasting
Traffic flow forecasting is a topic that has been studied for several decades Smith and Demetsky (1997); Sun et al. (2006); Yang et al. (2016); Song et al. (2020); Li et al. (2024c, a). Early efforts mainly focus on traditional models Smith and Demetsky (1997); Asadi et al. (2012). With the success of deep learning, deep neural networks have become mainstream in this field. One line of research adopts recurrent neural networks (RNNs) Hochreiter (1997) and graph neural networks (GNNs) Kipf and Welling (2016), where the GNNs and RNNs capture the spatial and temporal relations, respectively Li et al. (2018); Wang et al. (2020); Chen et al. (2022b); Li et al. (2023); Weng et al. (2023).
To jointly model spatial and temporal relations, another line of research utilizes GNNs in both dimensions Song et al. (2020); Li and Zhu (2021); Lan et al. (2022); Chen et al. (2024a). As simple graphs only capture pair-wise relations, some works take a step further, introducing hypergraph neural networks (HGNNs) to capture non-pair-wise spatio-temporal relations Luo et al. (2022); Wang and Zhu (2022); Sun et al. (2022); Zhao et al. (2023); Wang et al. (2024). In this work, we take advantage of the benefits from both sides, designing a dual-branch architecture, which better captures the rich structures of spatio-temporal relations.
Figure 2: The framework of the proposed LEAF, consisting of a dual-branch predictor and an LLM-based selector. The predictor generates forecasts of traffic flows with the graph and hypergraph branches. The selector constructs a choice set and selects the best option using an LLM. The selection results are used to supervise the predictor.
@@ 2.2 Large Language Models
In recent years, large language models have drawn increased attention, both within the community of natural language processing Chen et al. (2024b); Yin et al. (2024); Luo et al. (2025) and beyond, including healthcare Liévin et al. (2024); Van Veen et al. (2024); Labrak et al. (2024), education Milano et al. (2023), legal technology Lai et al. (2024), economics Li et al. (2024b), recommendation Chen et al. (2024b); Bao et al. (2024), code understanding Zhao et al. (2025b), and transportation Liu et al. (2024a); Guo et al. (2024b); Ren et al. (2024).
Among these applications, traffic flow forecasting is an important one, as it serves as the foundation of many downstream tasks in the field of intelligent transportation systems Boukerche et al. (2020). The success of LLMs in language has inspired a plethora of works in this task. Some works adopt the architecture of LLMs for building transformer-based traffic flow predictors Cai et al. (2020); Xu et al. (2020); Chen et al. (2022a); Liu et al. (2023); Jiang et al. (2023); Zou et al. (2024). However, they typically require a large amount of data for training.
Another line of research attempts to equip LLMs with the ability to predict future traffic flows based on the history and specific situations Zheng et al. (2023b); Guo et al. (2024a); Li et al. (2024c); Yuan et al. (2024); Han et al. (2024). Although LLMs have shown promising results in understanding time series Yu et al. (2023); Jin et al. (2023); Gruver et al. (2024); Koval et al. (2024); Gruver et al. (2024); Ansari et al. (2024); Liu et al. (2024b); Woo et al. (2024) or temporal events Xia et al. (2024); Hu et al. (2024), the traffic data involves complex spatio-temporal relations challenging LLMs’ generative ability. In this work, we show that one can instead utilize their discriminative ability to enhance existing traffic flow forecasting models.
@@ 3 Methodology
Problem Setup. We follow the standard setup for traffic flow forecasting Song et al. (2020), where there is a road network denoted as $\mathcal{G}=\langle\mathcal{V},\mathcal{E},\bm{A}\rangle$. $\mathcal{V}$ denotes the set of $N$ vertices (i.e. the sensors in the city), $\mathcal{E}$ denotes the set of edges, and $\bm{A}\in\mathbb{R}^{N\times N}$ represents the adjacency matrix. In this road network, historical traffic flows can be represented as $\bm{X}=(\bm{X}_{1},\bm{X}_{2},\cdots,\bm{X}_{T})\in\mathbb{R}^{T\times N\times F}$, in which $T$ is the length of historical observation and $F$ is the dimension of input features. The goal is to predict the future of traffic flows with the length of $T^{\prime}$, denoted as $\bm{X}^{\prime}=(\bm{X}^{\prime}_{1},\bm{X}^{\prime}_{2},\cdots,\bm{X}^{\prime}_{T^{\prime}})\in\mathbb{R}^{T^{\prime}\times N\times F}$.
Framework Overview. To solve the problem with the help of LLMs, we propose a novel framework termed Large Language Model Enhanced Traffic Flow Predictor (LEAF), whose overview is illustrated in Figure 2. Specifically, to achieve LLM-enhanced prediction (Section 3.3), we design a dual-branch traffic flow predictor (Section 3.1) and an LLM-based selector (Section 3.2). The predictor consists of a graph neural network branch capturing pair-wise spatio-temporal relations, and a hypergraph neural network branch capturing non-pair-wise relations. During training, the predictor is first pretrained using the training data. During test time, we apply transformations to the forecasting results of the predictor to obtain a variety of choices, among which the selector chooses the best one with a frozen LLM. The selection results are then used to fine-tune the dual-branch predictor.
@@ 3.1 Dual-branch Traffic Flow Predictor
Previous works on traffic flow forecasting adopt the graph perspective Song et al. (2020); Zheng et al. (2023a) or the hypergraph perspective Wang et al. (2022); Zhao et al. (2023). The graph perspective propagates messages between pairs of nodes, which makes them adept in modeling pair-wise spatio-temporal relations (e.g. an accident affects adjacent locations). On the other hand, the hypergraph perspective propagates messages among groups of nodes, making them proficient in modeling non-pair-wise spatio-temporal relations (e.g. people move from the residential area to the business area). For LEAF, we aim to take the benefits from both sides by constructing a dual-branch predictor and letting the LLM select.
@@
where $\sigma(\cdot)$ is the activation layer and $\bm{W}_{G}^{(l)}\in\mathbb{R}^{d\times d}$ is the weight matrix. $\widehat{\bm{A}^{ST}}=\bm{D}^{-1/2}\bm{A}^{ST}\bm{D}^{-1/2}$ denotes the normalized version of the adjacency matrix, where $\bm{D}$ is the degree matrix Kipf and Welling (2016); Song et al. (2020). By adopting graph convolutions in Eq. 3, the information from one spatio-temporal vertex can be propagated to its neighbors as defined in $\mathcal{E}^{ST}$ in Eq. 3.1, and thus this branch models pair-wise spatio-temporal relations.
The Hypergraph Branch. Although the graph branch is adept at capturing pair-wise relations, the complex traffic patterns contain non-pair-wise relations. For example, in the morning rush hours, people move from the residential area (which is a set of vertices in a hyperedge) to the business area (which is another hyperedge). The vertices in one hyperedge share common patterns and the hyperedges affect each other. To model non-pair-wise relations, hypergraphs are adopted. For a hypergraph, its incidence matrix $\bm{I}_{H}\in\mathbb{R}^{NT\times m}$ describes the assignment of $NT$ vertices to $m$ hyperedges. As the incidence matrix is not given as input, we resort to a learnable one with low-rank decomposition Zhao et al. (2023):
@@
$\bm{X}^{(l)}_{H}=\bm{I}_{H}\left(\bm{I}_{H}^{\top}\bm{X}^{(l-1)}_{H}+\sigma\left(\bm{W}_{E}\bm{I}_{H}^{\top}\bm{X}^{(l-1)}_{H}\right)\right),$ (5)
@@
3.2 LLM-Based Selector
In Section 3.1, we obtain different prediction results from different branches, denoted as $\bm{Y}^{G}$ and $\bm{Y}^{H}$. Most previous works Li et al. (2024c); Ren et al. (2024); Liang et al. (2024) directly generate these predictions, which is challenging since the complex spatio-temporal relations are hard to express in texts. By comparison, the LLM-based selector aims to choose the best prediction, using the internal knowledge of a frozen LLM and the constructed prompts. As the traffic networks are usually large, we break the result $\bm{Y}\in\{\bm{Y}^{G},\bm{Y}^{H}\}\subset\mathbb{R}^{T^{\prime}\times N}$ for individual vertices, i.e. $\bm{Y}=[\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{N}]$, where $\bm{y}_{i}\in\mathbb{R}^{T^{\prime}}$.
Choice Set Construction. In practice, we want to give the LLM-based selector more choices so that it has the potential to make better predictions. Therefore, we introduce several transformations: smoothing, upward trend, downward trend, overestimating, and underestimating. The set of transformations is denoted as $\mathcal{T}$. For vertex $i\in\mathcal{V}$, the choice set is determined as follows: